A man invests $2000 in a bank paying 5% simple interest per annum and $3000 in another bank paying 6% simple interest per annum. What rate percent does he recieve on the investments altogether?

Found 2 solutions by mathispowerful, andkon:
Answer by mathispowerful(115)


Rate equals interest divided by the principle in this case
Interest = 2000*5% + 3000*6% = $280
the principle is 2000 + 3000 = 5000
so rate R = 280/5000 = 0.056 = 5.6%
so the answer is 5.6%

Answer by andkon(15)